Can you please help me solve the problem? It's solving radical equations.
2x + square root (x+1) =8
The only thing in the square root is x + 1. I wasn't sure what sign to use on the keyboard for that.
I know the answer is 3, but I don't understand how to get it. I know you have to square both sides. Thank you.

Answer by stanbon(75887)

2x + square root (x+1) =8
--------------------------
Isolate the square root term:
sqrt(x+1) = 8-2x
---
Square both sides to get:
x+1 = 64 -32x + 4x^2
---
Modify:
4x^2 -33x +63 = 0
(x-3)(4x-21) = 0
x = 3 or x = 21/4
------------------
Checking those "answers":
Try x = 3 in 2x + sqrt (x+1) =8 and you get: 6 + sqrt(4) = 8
6 + 2 = 8
Good answer.
---------------------
Try x = 21/4 in 2x + sqrt(x+1) =8 and you get
(21/2) + sqrt(25/4) = 8
(21/2) + 5/2 = 8
13 = 8
That is false so x=21/4 is an extraneous answer.
====================================================
Cheers,
Stan H.